OpenStudy (anonymous):

A ramp has an angle of elevation of 19 and a length of 35 ft. What is the height in inches of the ramp at its high end?

OpenStudy (anonymous):

the angle is 19 degree right?

OpenStudy (anonymous):

yes the angle of the degree is 19

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sin 19 = x/35 therefore 35 sin 19 = x = 5.24570235 i think

OpenStudy (anonymous):

sin x = opp/hyp

OpenStudy (anonymous):

yep thts right

OpenStudy (anonymous):

x=11.40

OpenStudy (anonymous):

the height is 11.40

OpenStudy (anonymous):

in ft

OpenStudy (anonymous):

136.8in

OpenStudy (anonymous):

anything else u need to ask

OpenStudy (anonymous):

how did you get 11.4? :/

OpenStudy (anonymous):

11.4 ft 136.8in as the question asked for

OpenStudy (anonymous):

I think that is too big to be the answer. Think about it. how could elevation of only 19 degrees give you 11 ft in 35 ft of length? :/ It should be ~5

OpenStudy (anonymous):

but the question is in ft

OpenStudy (anonymous):

its 35 ft in lenght and 11.4 ft in height